OpenStudy (aa1996):

The diameter of a water tower tank is 10ft , and its height is 9ft . What is the volume of water that can be stored in the tank? (Use the value3.14 for π , and round your answer to the nearest cubic foot. Be sure to include the correct unit in your answer.) can you help me please

OpenStudy (anonymous):

\[ \text{As you know, }A=\pi r^2\text{ gives the area of a circle of radius }r.\\ \text{Similarly, }V=\pi r^2h\text{ gives the volume of a cylinder of radius }r\\ \text{and height }h.\\\\ \text{We're given that...}\\ \ \ \ r=\frac{10\text{ ft}}2=5\text{ ft}\\ \ \ \ h=9\text{ ft}\\ \ \ \ \pi=3.14\\ \text{So, all we must do now is merely plug in.}\\ \ \ \ V=\pi r^2h\\ \ \ \ \ \ \ = 3.14\times5\text{ ft}\times9\text{ ft}\\ \ \ \ \ \ \ = 141.3\text{ ft}^3\\\ \ \ \ \ \ \simeq 141\text{ ft}^3 \]

OpenStudy (aa1996):

thank you sooo much i understand it now =)